# Quantization-Free Autoregressive Action Transformer

### v1

@@ 4.1 Baselines and Performance Metrics
PushT Final IoU (/1$\cdot$/1⋅ / 1) 0.39 0.73 0.74 0.78 0.80
@@
Kitchen Goals (/4$\cdot$/4⋅ / 4) 3.07 2.62 3.44 3.66 3.84
@@
UR3 BlockPush Goals (/2$\cdot$/2⋅ / 2) 1.59 1.83 1.82 1.84 1.96
@@
Conditional Kitchen Goals (/4$\cdot$/4⋅ / 4) 0.15 3.75 3.47 3.78 3.78
Conditional UR3 BlockPush Goals (/2$\cdot$/2⋅ / 2) 0.00 1.14 0.92 1.94 1.96
@@
We compare Q-FAT against a range of state-of-the-art baselines. We restrict our focus to comparing against diffusion and transformer-based policy learning methods. For unconditional tasks, the baselines include: (1) BeT (Shafiullah et al., 2022), which performs action discretization using k-means clustering and models the action distribution with a transformer decoder; (2) VQ-BeT (Lee et al., 2024), an extension of BeT that leverages vector quantization via a hierarchical Variational Autoencoder (VQ-VAE) (Zeghidour et al., 2021) for action discretization; and (3) two variants of Diffusion Policy (Chi et al., 2023), namely a convolutional-based implementation and a transformer-based implementation. For goal-conditional tasks, we extend the comparison to include conditional versions of BeT and VQ-BeT, as well as two additional baselines: BESO (Reuss et al., 2023), a denoising diffusion-based method with a conditional variant (C-BESO) and a classifier-free guided variant (CFG-BESO).
@@ 4.2 Quality and Diversity of Q-FAT’s Actions
We assess performance using task success rate and induced behavioral entropy. Task success is measured differently per environment: in Kitchen, Multimodal Ant and UR3 BlockPush, the task success rate corresponds to the number of goals achieved per roll-out, while in PushT, it is quantified using the final Intersection over Union (IoU) between the object’s position and the target area. Behavioral diversity is measured by computing task completion sequence entropy (length 4 for Kitchen and Multimodal Ant, length 2 for UR3 BlockPush), reflecting variability in task completion. Baseline performance follows metrics from Lee et al. (2024), and results are reported as the average success rate of the best checkpoint over 1000 episodes, including behavioral entropy where applicable.
@@
Q-FAT achieves state-of-the-art performance on the success metrics, as shown in Table 1. The performance gap observed with BeT can be attributed to the limitations of k-means clustering in high-dimensional spaces, which may hinder its ability to effectively capture complex action distributions. This is particularly exacerbated due to BeT attempting to discretize the whole unconditional action space, and not the conditional one. In the case of VQ-BeT, the sequential nature of first predicting a discrete action token followed by an action correction step introduces two potential sources of error: inaccuracies in token prediction and errors in the subsequent correction mechanism. This places significant reliance on the vector quantization network, which is challenging to optimize due to the need to discretize the entire action space and the two-stage training process required—first for the quantization network and then for the policy. For diffusion-based policies, the absence of an explicit likelihood function makes it difficult to control the variance of sampled actions, particularly in the presence of noisy data. This can be particularly problematic in fine-grained control tasks, where precise action execution is crucial for success. Consequently, while these approaches demonstrate strong performance in certain scenarios, they may struggle in tasks requiring high accuracy and reliability under noisy data. As for assessing the behavioral diversity, Q-FAT largely outperforms baselines on the behavioral entropy in the Kitchen and the Multimodal Ant environments, while marginally under-performing in UR3 BlockPush compared to BeT, VQ-BeT and the transformer-based Diffusion Policy (Figure  6).
@@ 4.3 Effects of Sampling Techniques
We assessed the effect of the different sampling techniques discussed in Section 3.2. We found that down-scaling the component variances had a 7% and a 6.5% increase in the success performance metrics in the Kitchen environment and PushT, respectively, compared to the GMM sampling, while only seeing a marginal 2% drop in the diversity of the generated behaviors in the Kitchen environment. The modes sampling matched the performance of the variance down-scaling on the evaluation metrics. This could be attributed to the fact that the learned mixture components are sufficiently distant in the environments we explored, thus the effects discussed in Section 3.2 are not pronounced. For UR3 BlockPush, the sampling methods performed similarly on the performance metrics. In the Multimodal Ant environment, we observed a substantial 65% decrease in performance when using variance downscaling and mode sampling at $k=2$, compared to direct sampling from the GMM. Notably, this significant performance drop was only observed at $k=2$; at higher values of $k$, performance differences between all sampling methods became negligible. We expect the performance gap between the variance reduction sampling and the modes sampling to become more severe in higher-dimensional action spaces, but leave this exploration for future work.
Despite the quantitative metrics not showing significant differences between the variance down-scaling method and the mode sampling, we visualize the trajectories produced by both sampling techniques on a toy multi-route environment introduced in Shafiullah et al. (2022) and see the effects of the irreducible variance in Figure 5. We further display a similar effect in the PushT environment in Figure 7.
@@ 4.4 Effect of the Number of Mixtures
We systematically evaluated the effect of varying the number of mixture components $k$ in Q-FAT, where $k$ controls the number of modes in the conditional action distribution for a given state context. Specifically, we considered $k\in\{2,4,8,16\}$ in two benchmark domains, the Kitchen and UR3 BlockPush environments. Our results indicate that Q-FAT’s success performance is highly robust across these values of $k$, with negligible degradation in performance. In the Kitchen environment, we observed an 18% increase in behavioral entropy when $k$ was increased from 2 to 4, after which entropy remained approximately constant as $k$ continued to increase. In contrast, in the UR3 BlockPush environment, the behavioral entropy remained stable regardless of $k$.
@@
The key to achieving this robustness is careful initialization. In particular, we found it important to initialize each mixture component’s mean to be far from the others. Concretely, we placed these means on a hyper-cube spanning the range $[-1,1]$ per dimension, maximizing their mutual separation, and assigned each component a unit variance. The range $[-1,1]$ is a reasonable choice, as actions are typically bounded in practice and normalized in our experiments.
@@
To investigate how the mixture components were utilized during inference, we tracked the average number of active components over the course of training (see Figure 8). We observed that the model effectively prunes redundant mixture components as training progresses, with the pruning effect becoming more pronounced when more components are initially employed. Interestingly, in the Kitchen environment, which has a nine dimensional action space, less pruning occurs as $k$ increases compared to the UR3 environment, whose action space is only two dimensional.
@@
A plausible explanation for this discrepancy lies in the dimensionality of the action space. In higher dimensions, each Gaussian component covers a proportionally smaller volume of the action space. Consequently, the model requires more components to adequately represent the conditional action distribution, and therefore prunes fewer mixture components. In lower-dimensional environments, by contrast, each component can cover a larger fraction of the data, which makes aggressive pruning more viable.
@@
We further plotted the distribution over the number of active modes for the Kitchen and the UR3 BlockPush environments. We observed that approximately 70% of the course of a trajectory is uni-modal (Figure 9). This shows that multi-modality is required only a minority of the time during a rollout. This observation could be useful for future work to exploit this uni-modality computationally.
Figure 8: We observe that Q-FAT has a pruning effect, reducing the number of active mixtures at inference as training proceeds. We define the active number of mixtures at each training step as the count of mixture components whose probability is at least $10^{-2}$ times the maximum mixture probability at each inference step. We averaged the number of active mixtures over the rollout horizon for each environment and display the ratio of active mixtures for each $k$ value.
Figure 9: We analyzed the mode distribution over 100 rollouts for trained models with $k=4$ in the Kitchen and UR3 environments. Our findings indicate that in both environments, approximately 70% of the trajectory maintains only a single active mode, suggesting that the policy remains largely unimodal throughout each rollout.
4.5 Training Details
For our experiments, we used the minGPT (Brown et al., 2020) backbone as the decoder-only transformer implementation. We added the hyperparameters used for each of the environments in Table 2.
During training, we observed instances where both validation and training losses decreased, yet the environment rewards declined. We attribute this phenomenon to causal confusion (De Haan et al., 2019), where the model overfitslations in the recent sequence of states. Instead of learning true causal relationships, it relies on temporally correlated but non-causal patterns to predict future actions, leading to low rewards. To mitigate this issue, we introduced history masking, wherein ein whichment in the context (except for the current state) is masked with a certain probability. We applied this technique in both the Kitchen and Multimodal Ant environments (with probabilities 0.7 and 0.3 respectively), which led to an improved correlation between environment rewards and validation loss.
For the Image PushT, we fine-tuned a ResNet18 encoder (He et al., 2016) by extracting the feature maps from the first four layers in order to construct low level features (Parisi et al., 2022), which are then spatially average-pooled and fed into the transformer with a dimensionality of 1024. We further applied data augmentation, namely color jitter, random cropping, and random gray-scaling of the image with a probability of 0.5. We found this to be important to preserve the invariance of the pre-trained ResNet and not overfit to our dataset.
In the goal conditional tasks, we used a sequence of future states that has the same length as the state history. At inference time, we randomly sample a trajectory from the dataset and track it, computing the reward only if the agent achieved the goal in the same order that was achieved by the reference trajectory.
@@ 5 Related Work
Learning to act from offline data has received significant attention over the years due to its vast potential in practical applications. This line of research can be divided into two categories: offline reinforcement learning (Lange et al., 2012; Levine et al., 2020; Kumar et al., 2020; Chen et al., 2021) and imitation learning (Pari et al., 2021; Peng et al., 2018, 2021; Ho & Ermon, 2016). In offline reinforcement learning, on the one hand, the goal is to learn a policy from demonstrations with suboptimal behaviors, which necessitates the access to rewards in the demonstration dataset. Imitation learning, on the other hand, assumes that the trajectories are collected from near optimal behaviors, avoiding access to reward signals in the collected datasets. This increases the attractiveness of imitation learning in practical setups, since reward signals are hard to define (Devidze et al., 2021; Kumar et al., 2022). Q-FAT falls under the framework of behavioral cloning, which is a form of imitation learning where the policy learning is treated as a supervised learning problem.
@@
Complex generative models have been employed to capture the full data diversity due to the complex and multimodal behavior in human and robot collected datasets. Earlier work has explored the use of Gaussian processes (Shon et al., 2005), energy-based models (Florence et al., 2022) and generative adversarial networks (Kuefler et al., 2017) as policy parametrizations. More recently, more attention has been directed towards more powerful models, such as transformers (Shafiullah et al., 2022; Lee et al., 2024; Mete et al., 2024; Wu et al., 2023) and diffusion models (Chi et al., 2023; Wang et al., 2022; Psenka et al., 2023; Pearce et al., 2023). Unlike other transformer-based policies, Q-FAT uses a continuous parametrization of the output by predicting a GMM. Note that this idea has been explored before using LSTM backbones (Mandlekar et al., 2021; Shafiullah et al., 2022), however, it has been shown to produce poor performance (Shafiullah et al., 2022; Lee et al., 2024), which we attribute to the instability of training LSTM networks and their limited capacity to model long sequences (Hochreiter, 1997).
@@ 6 Discussion and Future Work
In this work, we introduced Q-FAT, a quantization-free action transformer that achieves state-of-the-art performance across various simulated robotics environments, opening new research directions in both behavioral modeling and reinforcement learning. Compared to prior work  (Lee et al., 2024), Q-FAT overcomes the challenging non-differentiable step in action discretization while maintaining or improving performance.
@@
Impact Statement
@@
Acknowledgment
We would like to thank Youssef Saied for the valuable discussions throughout the project and Nicolas Nguyen for reviewing the document. C.V. and Z.S. are funded by the German Research Foundation (DFG) under both the project 468806714 of the Emmy Noether Programme and under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. M.M. is funded by the German Research Foundation (DFG) under the project 456587626 of the Emmy Noether Programme. We also thank the international Max Planck Research School for Intelligent Systems (IMPRS-IS) for the support.
@@ Appendix A GMM Properties
where $\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ is the multivariate Gaussian distribution for the $i$-th component. The mean and the covariance of $\mathbf{X}$ are then given by:
@@ Appendix B GMM Mode Finding and Sampling
and a critical point $\mathbf{x}^{*}$ is a mode if: $\nabla p(\mathbf{x}^{*})=\mathbf{0}\text{ and }\nabla^{2}p(\mathbf{x}^{*})\text{ is negative definite.}$.
While GMMs are often parameterized by a finite number of components, the resulting distribution can exhibit a richer structure, particularly in higher-dimensional spaces. Specifically, in dimensions greater than two ($d>2$), a GMM can possess more modes (local maxima) than the number of its constituent components (Carreira-Perpinan, 2000). This phenomenon arises due to the interplay between the covariance structures and the relative positions of the Gaussian components. In higher dimensions, the overlapping regions of different components can create multiple peaks in the probability density function that cannot be attributed to individual components.
@@
The fixed-point iteration has the following interpretation: each iteration shifts the current iterate after $t$ iterations closer to the modes (high-density regions) of the underlying density defined by the mixture of Gaussians. In practice, it has been shown that initializing $\mathbf{x}^{(0)}$ with the component means $\boldsymbol{\mu}_{i}$ results in fast convergence (Carreira-Perpinan, 2000), which we also observe in our experiments. To account for the fact that the number of modes could exceed the number mixture components, we add extra initialization points from the convex-hull of the component centroids capturing the most important modes.
@@
Since each $\boldsymbol{\Sigma}_{i}^{-1}$ is also diagonal (with entries $\frac{1}{\sigma_{i,j}^{2}}$ along the diagonal), the vector and matrix operations inside $T(\mathbf{x})$ decouple across dimensions. Hence the inverse of $\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\,\boldsymbol{\Sigma}_{i}^{-1}$ reduces to a diagonal matrix whose components can be computed efficiently. As a result, each coordinate of the updated point can be determined independently, making the mean-shift iteration particularly fast on modern hardware.
@@
To sample the modes proportional to their coverage of the data support, we approximate the integral of the GMM density in the local neighborhood of each mode using a Laplace approximation and use the value of the integral around each mode to compute the mode’s relative weight. We summarize the full mode-finding and sampling procedure in Algorithm 1.
@@
4:  for $x\in\mathcal{X}$ do
5:     Iteration counter $i\leftarrow 0$
6:     Define the operator $T(.)$ as in Equation 16
7:     repeat
8:        Update $\mathbf{x}$ via fixed-point iteration: $\mathbf{x}\leftarrow T(\mathbf{x})$
9:        Increment $i$
@@
11:     Compute $\mathbf{H}=\nabla^{2}p(\mathbf{x})$
12:     if $\max(\text{eigenvalues}(\mathbf{H}))<0$ then
13:        Merge nearby modes within a distance threshold, keeping the highest-probability mode
14:        Add mode to $\mathcal{M}$
15:     end if
16:  end for
17:
18:  Compute Mode Weights: $w_{j}=p(\mathbf{x}_{j})\bigl{|}\!-\!\nabla^{2}\log p(\mathbf{x}_{j})\bigr{|}^{-1/2},\,j=0,\dots,|\mathcal{M}|$
19:  Sample mode: Sample $j\sim\text{Categorical}\left(\{w_{j}\}_{j=0}^{|\mathcal{M}|}\right)$
20:  Return: $\mathbf{x}_{j}$

### v3

@@ 4.1 Baselines and Performance Metrics
PushT Final IoU ($\cdot/1$) 0.39 0.73 0.74 0.78 0.80
@@
Kitchen Goals ($\cdot/4$) 3.07 2.62 3.44 3.66 3.84
Image Kitchen 2.48 3.11 3.01 2.98 3.15
@@
UR3 BlockPush Goals ($\cdot/2$) 1.59 1.83 1.82 1.84 1.99
BlockPush 1.67 0.47 1.93 1.79 1.77
@@
Cond. Kitchen Goals ($\cdot/4$) 0.15 3.75 3.47 3.78 3.78
Cond. UR3 BlockPush Goals ($\cdot/2$) 0.00 1.14 0.92 1.94 1.96
@@
We compare Q-FAT against a range of state-of-the-art baselines. We restrict our focus to comparing against diffusion and transformer-based policy learning methods. For unconditional tasks, the baselines include: (1) BeT [45], which performs action discretization using k-means clustering and models the action distribution with a transformer decoder; (2) VQ-BeT [26], an extension of BeT that leverages vector quantization via a hierarchical Variational Autoencoder (VQ-VAE) [57] for action discretization; and (3) two variants of Diffusion Policy [7], namely a convolutional-based implementation and a transformer-based implementation. For goal-conditional tasks, we extend the comparison to include conditional versions of BeT and VQ-BeT, as well as two additional baselines: BESO [42], a denoising diffusion-based method with a conditional variant (C-BESO) and a classifier-free guided variant (CFG-BESO).
@@ 4.2 Quality and Diversity of Q-FAT’s Actions
We assess performance using task success rate and induced behavioral entropy. Task success is measured differently per environment: in Kitchen, Multimodal Ant and UR3 BlockPush, the task success rate corresponds to the number of goals achieved per roll-out, while in PushT, task success is quantified using the final Intersection over Union (IoU) between the object’s position and the target area. Behavioral diversity is measured by computing task completion sequence entropy (length 4 for Kitchen and Multimodal Ant, length 2 for UR3 BlockPush), reflecting variability in task completion. Baseline performance follows metrics from Lee et al. [26], and results are reported as the average success rate of the best checkpoint over 1000 episodes, including behavioral entropy where applicable.
@@
Q-FAT achieves state-of-the-art performance on the success metrics, as shown in Table 1. The performance gap observed with BeT can be attributed to the limitations of k-means clustering in high-dimensional spaces, which may hinder its ability to effectively capture complex action distributions. This is particularly exacerbated due to BeT attempting to discretize the whole unconditional action space, and not the conditional one. In the case of VQ-BeT, the sequential nature of first predicting a discrete action token followed by an action correction step introduces two potential sources of error: inaccuracies in token prediction and errors in the subsequent correction mechanism. This places significant reliance on the vector quantization network, which is challenging to optimize due to the need to discretize the entire action space and the required two-stage training process — first for the quantization network and then for the policy. For diffusion-based policies, the absence of an explicit likelihood function makes it difficult to control the variance of sampled actions, particularly in the presence of noisy data. This can be particularly problematic in fine-grained control tasks, where precise action execution is crucial for success. Consequently, while these approaches demonstrate strong performance in certain scenarios, they may struggle in tasks requiring high accuracy and reliability under noisy data. Q-FAT largely outperforms or matches the behavioral entropy of all baselines (Figure  3), which assesses diversity across the generated motion.
@@ 4.3 Effects of Sampling Techniques
We assessed the effect of the different sampling techniques discussed in Section 3.2.
(a) Dataset
(b) GMM Sampling
(c) $\sigma$-scaled Sampling
(d) Mode Sampling
Figure 4: Sampling techniques from an 8-mixture Q-FAT policy on a multiroute environment [45]. (a) Raw dataset with two pairs of equally likely paths (blue and red) from start to target (green). (b) Direct GMM sampling yields noisy samples due to the captured dataset noise. (c) Variance scaling ($10^{-8}$)reduces per-component variance but not the variance from inter-component distances.(d) Mode sampling largely suppresses noise while preserving dataset multimodality.
We found that down-scaling the component variances had a 7% and a 6.5% increase in the success performance metrics in the Kitchen environment and PushT, respectively, compared to the GMM sampling, while only seeing a marginal 2% drop in the diversity of the generated behaviors in the Kitchen environment.
The modes sampling matched the performance of the variance down-scaling on the evaluation metrics. This could be attributed to the fact that the learned mixture components are sufficiently distant in the environments we explored, thus the effects discussed in Section 3.2 are not pronounced.
Despite the quantitative metrics not showing significant differences between the variance down-scaling method and the mode sampling, we visualize the trajectories produced by both sampling techniques on a toy multi-route environment introduced by Shafiullah et al. [45] and see the effects of the irreducible variance in Figure 4. We further display a similar effect in the PushT environment in Figure 5.
@@ 4.4 Effect of the Number of Mixtures
Figure 6: The left two plots display the evolution of the number of active mixtures during inference as training progresses, showing a clear pruning effect. On the rightmost plot, we display the distribution of the number of active modes at each timestep during inference for a converged model. One can see that trajectories remain largely unimodal.
We systematically evaluated the effect of varying the number of mixture components $k$ in Q-FAT, where $k$ controls the number of modes in the conditional action distribution for a given state context.
@@
Specifically, we considered $k\in\{2,4,8,16\}$ in two benchmark domains, the Kitchen and UR3 BlockPush environments. Our results indicate that Q-FAT’s success performance is highly robust across these values of $k$, with negligible degradation in performance. In the Kitchen environment, we observed an 18% increase in behavioral entropy when $k$ was increased from 2 to 4, after which entropy remained approximately constant as $k$ continued to increase. In contrast, in the UR3 BlockPush environment, the behavioral entropy remained stable regardless of $k$. The key to achieving this robustness is careful initialization.
@@
In particular, we found it important to initialize each mixture component’s mean to be distant from the others. Concretely, we placed these means on a hyper-cube spanning the range [1,1][-1,1] per dimension, maximizing their mutual separation, and assigned each component a unit variance. The range [1,1][-1,1] is a reasonable choice, as actions are typically bounded in practice and normalized in our experiments.
@@
To investigate how the mixture components were utilized during inference, we tracked the average number of active components over the course of training (see Figure 6, left). We observed that the model effectively prunes redundant mixture components as training progresses, with the pruning effect becoming more pronounced when more components are initially employed. Interestingly, in the Kitchen environment, which has a nine dimensional action space, less pruning occurs as $k$ increases compared to the UR3 environment, whose action space is only two dimensional. A plausible explanation for this discrepancy lies in the dimensionality of the action space.
@@
In higher dimensions, each Gaussian component covers a proportionally smaller volume of the action space. Consequently, the model requires more components to adequately represent the conditional action distribution, and therefore prunes fewer mixture components. In lower-dimensional environments, by contrast, each component can cover a larger fraction of the data, which makes aggressive pruning more viable.
@@
We further plotted the distribution over the number of active modes for the Kitchen and the UR3 BlockPush environments. We observed that approximately 70% of the course of a trajectory is uni-modal (Figure 6, right). This shows that multi-modality is required only a minority of the time during a rollout. This observation could be useful for future work to exploit this uni-modality computationally.
4.5 Inference Time Analysis
To compare computational efficiency, we evaluated the inference times of Q-FAT and VQ-BeT on a 16 GB MacBook Pro CPU. The analysis utilized models trained in the Kitchen environment, with hyperparameters for Q-FAT specified in Table 3 and for VQ-BeT as described in [26]. Q-FAT is approximately 2x faster on the tested hardware due to its efficient action decoding head. The breakdown of inference time per model component is detailed in Table 2.
Table 2: Inference time breakdown for Q-FAT and VQ-BeT on the models used for the Kitchen environment. Q-FAT has a lighter action-decoding head, resulting in a signigicant speedup.
Component VQ-BeT (ms) Q-FAT (ms)
Shared Backbone (minGPT) $\sim$1.0 $\sim$1.0
Action Decoding Head $\sim$1.2 $\sim$0.1
Total Inference Time $\sim$2.2 $\sim$1.1
The primary performance difference arises from the action decoding head. While both models use an identical backbone architecture, VQ-BeT’s reliance on two separate 2-layer MLPs for action decoding creates a computational bottleneck. In contrast, Q-FAT’s use of a single, efficient linear action decoding layer results in a significant speedup.
@@ 5 Related Work
Learning to act from offline data has received significant attention over the years due to its vast potential in practical applications. This line of research can be divided into two categories: offline reinforcement learning [25, 27, 22, 6] and imitation learning [34, 37, 38, 13]. In offline reinforcement learning, on the one hand, the goal is to learn a policy from demonstrations with suboptimal behaviors, which necessitates the access to rewards in the demonstration dataset. Imitation learning, on the other hand, assumes that the trajectories are collected from near optimal behaviors, avoiding access to reward signals in the collected datasets. This increases the attractiveness of imitation learning in practical setups, since reward signals are hard to define [9, 23]. Q-FAT falls under the framework of behavioral cloning, which is a form of imitation learning where the policy learning is treated as a supervised learning problem.
@@
Complex generative models have been employed to capture the full data diversity due to the complex and multimodal behavior in human and robot datasets. Earlier work has explored the use of Gaussian processes [46], energy-based models [10] and generative adversarial networks [21] as policy parametrizations. More recently, more attention has been directed towards more powerful models, such as transformers [45, 26, 32, 55] and diffusion models [7, 54, 40, 36]. Unlike other transformer-based policies, Q-FAT uses a continuous parametrization of the output by predicting a GMM. Note that this idea has been explored before using LSTM backbones [29, 45], however, it has been shown to produce poor performance [45, 26], which we attribute to the instability of training LSTM networks and their limited capacity to model long sequences [14].
@@ 6 Discussion and Future Work
In this work, we introduced Q-FAT, a quantization-free action transformer that achieves state-of-the-art performance across various simulated robotics environments, opening new research directions in both behavioral modeling and reinforcement learning. Compared to prior work  [26], Q-FAT overcomes the challenging non-differentiable step in action discretization while maintaining or improving performance.
@@
7 Limitations
While we have demonstrated that Q-FAT is effective in learning complex action distributions, the underlying GMM loss function used to train the policy assumes a Euclidean action space. In some environments (e.g., humanoid robots), the action space has a Riemannian geometry, which could potentially result in difficulty during learning. This can be counteracted by mapping actions into a latent Euclidean space using an action autoencoder [19] and perform the GMM loss in the latent space. However, unlike previous work that uses discrete action encoders [26, 45], Q-FAT allows the flexibility of using continuous latent spaces, allowing end-to-end differentiability of the joint encoder and policy models. While this is an exciting research direction, we leave this for future work.
8 Impact Statement
@@
9 Acknowledgment
We would like to thank Youssef Saied for the valuable discussions throughout the project and Nicolas Nguyen for reviewing the manuscript. C.V. and Z.S. are funded by the German Research Foundation (DFG) under both the project 468806714 of the Emmy Noether Programme and under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. M.M. is funded by the German Research Foundation (DFG) under the project 456587626 of the Emmy Noether Programme. We also thank the international Max Planck Research School for Intelligent Systems (IMPRS-IS) for the support.
@@ Appendix A GMM Properties
where $\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ is the multivariate Gaussian distribution for the $i$-th component. The mean and the covariance of $\mathbf{X}$ are given by:
@@ Appendix B GMM Mode Finding and Sampling
and a critical point $\mathbf{x}^{*}$ is a mode if: $\nabla p(\mathbf{x}^{*})=\mathbf{0}\text{ and }\nabla^{2}p(\mathbf{x}^{*})\text{ is negative definite.}$
While GMMs are often parameterized by a finite number of components, the resulting distribution can exhibit a richer structure, particularly in higher-dimensional spaces. Specifically, in dimensions greater than two ($d>2$), a GMM can possess more modes (local maxima) than the number of its constituent components [5]. This phenomenon arises due to the interplay between the covariance structures and the relative positions of the Gaussian components. In higher dimensions, the overlapping regions of different components can create multiple peaks in the probability density function that cannot be attributed to individual components.
@@
The fixed-point iteration has the following interpretation: each iteration shifts the current iterate after $t$ iterations closer to the modes (high-density regions) of the underlying density defined by the mixture of Gaussians. In practice, it has been shown that initializing $\mathbf{x}^{(0)}$ with the component means $\boldsymbol{\mu}_{i}$ results in fast convergence [5], which we also observe in our experiments. To account for the fact that the number of modes could exceed the number mixture components, we add extra initialization points from the convex-hull of the component centroids capturing the most important modes.
@@
Since each $\boldsymbol{\Sigma}_{i}^{-1}$ is also diagonal (with entries $1/\sigma_{i,j}^{2}$ along the diagonal), the vector and matrix operations inside $T(\mathbf{x})$ decouple across dimensions. Hence the inverse of $\sum_{i=1}^{k}\pi_{i}\,\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\,\boldsymbol{\Sigma}_{i}^{-1}$ reduces to a diagonal matrix whose components can be computed efficiently. As a result, each coordinate of the updated point can be determined independently, making the mean-shift iteration particularly fast on modern hardware.
@@
To sample the modes proportional to their coverage of the data support, we approximate the integral of the GMM density in the local neighborhood of each mode using a Laplace approximation and use the value of the integral around each mode to compute the relative weight of the mode. We summarize the full mode-finding and sampling procedure in Algorithm 1.
@@
5:for $x\in\mathcal{X}$ do
6: Iteration counter $i\leftarrow 0$
7: Define the operator $T(.)$ as in Equation 16
8:repeat
9:  Update $\mathbf{x}$ via fixed-point iteration: $\mathbf{x}\leftarrow T(\mathbf{x})$
10:  Increment $i$
11:until $i\geq\text{max\_it}$ or $\|\mathbf{x}-\mathbf{x}_{\text{old}}\|<\epsilon$
12: Compute $\mathbf{H}=\nabla^{2}p(\mathbf{x})$
13:if $\max(\text{eigenvalues}(\mathbf{H}))<0$ then
14:  Merge nearby modes within a distance threshold, keeping the highest-probability mode
15:  Add mode to $\mathcal{M}$
16:end if
17:end for
18:
19:Compute Mode Weights:
@@
21:Sample mode: Sample $j\sim\text{Categorical}\left(\{w_{j}\}_{j=0}^{|\mathcal{M}|}\right)$
22:Return: $\mathbf{x}_{j}$